A student has two springs. Spring 1 is twice the mass of spring 2. The student first puts spring 1 into a massless pouch and suspends it from spring 2. Spring extension was noted to be 11.55cm. Then, they put spring 2 into a massless pouch and suspend it from spring 1. Spring extension was noted to be 16.16cm.

What is the ratio of k1/k2?
Answer #1

Let mass of spring-2 is m kg and mass of spring-1 is 2m kg

Extension of spring-2 due to mass of spring-1 = 11..55 cm

force constant k2 of spring-2 = 11.55 / 2m = ( 5.775 / m ) cm/N

Extension of spring-1 due to mass of spring-2 = 16.16 cm

force constant k1 of spring-1 = 16.16 / m = ( 16.16 / m ) cm/N

ratio of ( k1 / k2 ) = 16.16 / 5.775 2.8